(D-1)x + Dy = 0

Dx + 2Dy = 4"e^2t"


1
Expert's answer
2021-01-22T02:28:29-0500

Let us solve the system:


"\begin{cases}
(D-1)x + Dy = 0\\
Dx + 2Dy = 4e^{2t}
\end{cases}"


"\begin{cases}
Dy = -(D-1)x \\
Dx + 2Dy = 4e^{2t}
\end{cases}"


"\begin{cases}
Dy = -(D-1)x \\
Dx - 2(D-1)x = 4e^{2t}
\end{cases}"


"\begin{cases}
Dy = -(D-1)x \\
-(D-2)x = 4e^{2t}
\end{cases}"


"\begin{cases}
Dy = -(D-1)x \\
(D-2)xe^{-2t} = -4
\end{cases}"


"\begin{cases}
Dy = -(D-1)x \\
D(xe^{-2t}) = -4
\end{cases}"


"\begin{cases}
Dy = -(D-1)x \\
xe^{-2t} = -4t+C_1
\end{cases}"


"\begin{cases}
Dy = -(D-1)x \\
x = -4te^{2t}+C_1e^{2t}
\end{cases}"


"\begin{cases}
Dy = 4e^{2t}+8te^{2t}-2C_1e^{2t}-4te^{2t}+C_1e^{2t}= 4e^{2t}+4te^{2t}-C_1e^{2t}\\
x = -4te^{2t}+C_1e^{2t}
\end{cases}"


We conclude that the solution is of the following form:


"\begin{cases}
y =e^{2t}+2te^{2t}-\frac{1}{2}C_1e^{2t}+C_2\\
x = -4te^{2t}+C_1e^{2t}
\end{cases}"